# Inverse Problem Sampling in Latent Space Using Sequential Monte Carlo

### v1

@@ Abstract
In image processing, solving inverse problems is the task of finding plausible reconstructions of an image that was corrupted by some (usually known) degradation model. Commonly, this process is done using a generative image model that can guide the reconstruction towards solutions that appear natural. The success of diffusion models over the last few years has made them a leading candidate for this task. However, the sequential nature of diffusion models makes this conditional sampling process challenging. Furthermore, since diffusion models are often defined in the latent space of an autoencoder, the encoder-decoder transformations introduce additional difficulties. Here, we suggest a novel sampling method based on sequential Monte Carlo (SMC) in the latent space of diffusion models. We use the forward process of the diffusion model to add additional auxiliary observations and then perform an SMC sampling as part of the backward process. Empirical evaluations on ImageNet and FFHQ show the benefits of our approach over competing methods on various inverse problem tasks.
@@ 1 Introduction
Many important signal processing tasks can be viewed as inverse problems (Song et al., 2021b; Moliner et al., 2023; Daras et al., 2024; Chung et al., 2023b; Cardoso et al., 2023). In inverse problems, the objective is to obtain a clean signal ${\mathbf{x}}\in\mathbb{R}^{n}$ from a degraded observation ${\mathbf{y}}={\mathcal{A}}({\mathbf{x}})+\bm{\mathbf{\psi}}$, where ${\mathcal{A}}$ is usually a known irreversible mapping and $\bm{\mathbf{\psi}}$ is a Gaussian noise vector. Common applications that fit this framework include image deblurring, super-resolution, inpainting, and Gaussian denoising. The broad applicability of inverse problems makes them highly significant, as they encompass numerous real-world challenges, such as those found in digital image processing (Blackledge, 2005), wireless communication (Chen et al., 2021), seismology (Virieux & Operto, 2009), medical imaging (Song et al., 2021b; Chung et al., 2023c), and astronomy (Craig & Brown, 1986).
@@
Figure 1: LD-SMC solves inverse problems in the latent space by first augmenting the model with auxiliary observations using the forward diffusion process. Then, sampling from the posterior distribution $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ is done based on the backward diffusion process using sequential Monte Carlo. In the figure, ${\mathcal{D}}$ and ${\mathcal{A}}$ denote the decoder and the corruption operator respectively.
@@
Figure 2: Comparison between LD-SMC and baseline methods on inpainting of images from the ImageNet dataset.
@@
A simple solution proposed in (Chung et al., 2023b) is to approximate $p({\mathbf{y}}|{\mathbf{x}}_{t})$ with $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ for efficient sampling. Further work in (Wu et al., 2024) applied a sequential Monte Carlo (SMC) process to correct for this approximation. Another recent approach proposed in (Dou & Song, 2024) connected ${\mathbf{x}}_{t}$ to ${\mathbf{y}}$ by introducing a sequence of latent variables ${\mathbf{y}}_{t}$ and sampling sequentially from $p({\mathbf{x}}_{t:T}|{\mathbf{y}}_{t:T})$. While this approach has shown great potential, it is limited to linear corruption models. As such, it is not applicable for non-linear mappings ${\mathcal{A}}$ or common Latent Diffusion Models (LDMs) (Rombach et al., 2022) due to the nonlinearity of the decoder. This is a serious restriction, as many of the recent powerful and efficient models are LDMs (Esser et al., 2024).
@@
Both existing approaches have pros and cons. Using the $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ approximation can be helpful in capturing the large scale semantics of the image, but it might not be well suited for capturing the small details. On the other hand, the auxiliary ${\mathbf{y}}_{1:T}$ can aid in the finer details and could be used with LDMs. Here, we propose a method that combines these two approaches and strives to achieve the best of both worlds. Specifically, we augment the model with additional latent variables ${\mathbf{y}}_{t}$, one for each time step, and then apply posterior inference over the latent diffusion variables ${\mathbf{z}}_{t}$. To obtain a tractable sampling procedure, we derive a novel posterior approximation and define a new proposal distribution for the SMC sampling process (Doucet et al., 2001; Del Moral et al., 2012). Hence, we name our method Latent Diffusion Sequential Monte Carlo, or more concisely LD-SMC. Importantly, as our approach utilizes these auxiliary variables in the sampling process, we found that it was better suited to more challenging inverse problems such as inpainting. An illustration of our approach is shown in Figure 1.
We empirically validated our approach on the ImageNet (Russakovsky et al., 2015) and FFHQ (Karras et al., 2019) datasets. LD-SMC usually outperforms or is comparable to baseline methods on image deblurring and super-resolution tasks, and can significantly improve over baseline methods on inapainting tasks, especially on ImageNet which has more diversity in it.
To conclude, in this study we make the following contributions: (1) A novel SMC procedure for solving inverse problems using latent diffusion models; (2) A novel posterior approximation and proposal distribution to perform approximate posterior sampling; (3) LD-SMC outperforms baseline methods, especially on challenging inpainting tasks.
@@ 2 Background
Inverse Problems. In inverse problems one would like to recover a sample ${\mathbf{x}}\in{\mathbb{R}}^{n}$ from a corrupted version of it ${\mathbf{y}}\in{\mathbb{R}}^{m}$. Usually, the corruption model that acted on ${\mathbf{x}}$ is known, but the operation is irreversible (Tarantola, 2005). For instance, restoring a high-quality image from a low-quality one. We denote the corruption operator by ${\mathcal{A}}(\cdot)$, and assume that ${\mathbf{y}}={\mathcal{A}}({\mathbf{x}})+\bm{\mathbf{\psi}}$, where $\bm{\mathbf{\psi}}\sim\mathcal{N}(0,\tau^{2}{\mathbf{I}})$ has a known standard deviation $\tau$. In a more concise way, $p({\mathbf{y}}|{\mathbf{x}})=\mathcal{N}({\mathcal{A}}({\mathbf{x}}),\tau^{2}{\mathbf{I}})$. Common examples of inverse problems are inpainting, colorization, and deblurring. In general, solving inverse problem tasks is considered an ill-posed problem with many possible solutions ${\mathbf{x}}$ with equally high $p({\mathbf{y}}|{\mathbf{x}})$ values. Given a prior distribution $p({\mathbf{x}})$ over natural images, one standard approach to solving the inverse problem is to sample the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{y}}|{\mathbf{x}})p({\mathbf{x}})$.
Diffusion Models. Owing to their high-quality generation capabilities, in recent years diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have been leveraged as priors in inverse problems (Jalal et al., 2021; Song et al., 2021c). Here, we adopt the DDIM formulation (Song et al., 2021a) for the prior model, although our approach can work with other diffusion model formulations as well. Furthermore, since it is costly to apply the diffusion process in the pixel space, a common approach is to apply the diffusion model in the latent space given by an auto-encoder (Rombach et al., 2022). Applying diffusion models in the latent space allows us to sample high-quality images while reducing the computational resources needed by the model. Hence, designing models that effectively solve inverse problems using latent diffusion models is of great importance.
Denote by ${\mathbf{z}}_{1:T}$ the random variables in the latent space. Let $\alpha_{1:T}$, $\beta_{1:T}$ be the variance schedule of the diffusion process with $\beta_{t}\coloneqq 1-\alpha_{t}$. Also, denote by $\bar{\alpha}_{t}\coloneqq\prod_{j=1}^{t}\alpha_{j}$. Then, the DDIM sampling is done according to $p_{\theta}({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})=\mathcal{N}({\mathbf{z}}_{t-1}|\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t},t),\mathbf{\Sigma}(t))$, where $\theta$ are the parameters of the neural network and,
@@
$\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t},t)$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{{\mathbf{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\mathbf{\epsilon}}_{\theta}({\mathbf{z}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+$
$~{}~{}~{}~{}\sqrt{1-\bar{\alpha}_{t-1}-\sigma^{2}_{t}}\cdot\bm{\mathbf{\epsilon}}_{\theta}({\mathbf{z}}_{t},t).$
@@
As in (Dou & Song, 2024) we fix $\sigma_{t}=\eta\cdot\sqrt{\beta_{t}\cdot\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}$ with $\eta$ being a hyper-parameter. We denote the approximate posterior mean of $\mathbb{E}[{\mathbf{z}}_{0}|{\mathbf{z}}_{t}]$ by $\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t})\coloneqq\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\mathbf{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\mathbf{\epsilon}}_{\theta}({\mathbf{z}}_{t},t))$ (Robbins, 1956; Efron, 2011; Chung et al., 2023b).
@@
Sequential Monte Carlo (SMC). SMC is an important technique for sampling in probabilistic graphical models in which exact posterior inference is intractable. The SMC breaks the sampling process down to intermediate steps, allowing efficient sampling through a recursive procedure (Doucet et al., 2001; Del Moral et al., 2012; Naesseth et al., 2019; Chopin et al., 2020).
@@
One family of probabilistic models for which SMC is especially known is state-space models (SSMs), also known as Hidden Markov Models (HMMs). In general, the following quantities need to be defined in SSMs, (1) a prior distribution over the initial state $p({\mathbf{z}}_{T})$, (2) a transition distribution that defines the dynamics between states $p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})~{}\forall t<T$, and (3) a measurement model $p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})~{}\forall t<T$. The goal is to sample from the posterior distribution $p({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T-1})$. To do so, SMC starts by sampling $N$ particles $\{{\mathbf{z}}^{(i)}_{T}\}_{i=1}^{N}$ from the prior distribution. Then, at each step, given the previous particle set $\{{\mathbf{z}}^{(i)}_{t}\}_{i=1}^{N}$ new samples are taken from a proposal distribution ${\mathbf{z}}^{(i)}_{t-1}\sim\pi({\mathbf{z}}^{(i)}_{t-1}|{\mathbf{z}}^{(i)}_{t})~{}\forall i\in\{1,...,N\}$. The particles are then weighted and resampled according to the new proposed sequences $\{{\mathbf{z}}^{(i)}_{t-1:T}\}_{i=1}^{N}$. The proposal distribution serves as an approximation to the posterior distribution. Its support needs to contain the support of the posterior density. The weighting function corrects the approximation by assigning a weight to each particle to adjust its probability. And the resampling step aims to remove unlikely particles according to the model (Särkkä, 2013).
3 Related Work
Inverse problems have a long and evolving history, with methodologies that have undergone significant advances over the years (Daras et al., 2024). Recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021c) have emerged as effective priors for solving inverse problems in image data (Wang et al., 2023; Kawar et al., 2022; Chung et al., 2023b; Dou & Song, 2024; Rout et al., 2023; Song et al., 2023; Sun et al., 2024; Choi et al., 2021; Chung et al., 2023a).
In (Song et al., 2021c) it was shown that to sample from the posterior distribution, $p({\mathbf{x}}_{0}|{\mathbf{y}})$, one can solve a stochastic differential equation based on the prior score, $\nabla_{{\mathbf{x}}_{t}}\log~{}p_{\theta}({\mathbf{x}}_{t})$, and the conditional score, $\nabla_{{\mathbf{x}}_{t}}\log~{}p_{\theta}({\mathbf{y}}|{\mathbf{x}}_{t})$. Although the first term is easy to compute, the latter term requires integration over the full diffusion path from time $t$ to 00. A useful and easy-to-calculate approximation found in several studies is $p_{\theta}({\mathbf{y}}|{\mathbf{x}}_{t})\approx p_{\theta}({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$, which is readily available at each step (Chung et al., 2023b; Song et al., 2023; Wu et al., 2024). Specifically, Diffusion Posterior Sampling (DPS) (Chung et al., 2023b) uses this approximation for linear and non-linear inverse problems with Gaussian and Poisson likelihood models. $\Pi$GDM (Song et al., 2023) introduces pseudoinverse guidance by matching the denoising output and the corrupted image ${\mathbf{y}}$, via transformation of both through a ’pseudoinverse’ of the corruption model. DDNM (Wang et al., 2023) suggested to refine only the contents of the null space during the backward diffusion process. As such it is suited only for linear inverse problems. An additional category of inverse problem approaches that use diffusion models is designed with the objective of asymptotic exactness (Cardoso et al., 2023; Trippe et al., 2023; Wu et al., 2024; Dou & Song, 2024). These methods utilize SMC techniques targeting exact sampling from the posterior distribution $p({\mathbf{x}}_{0}|{\mathbf{y}})$. Specifically, SMC-Diff (Trippe et al., 2023) applies particle filtering for inpainting in motif scaffolding, and MCGDiff (Cardoso et al., 2023) is designed for linear inverse problems only. Hence, both approaches are not suited for inverse problems with latent-space diffusion models. TDS (Wu et al., 2024), a recent SMC-based method, solves general inverse problem tasks using the twisting technique. This method also uses the approximation of DPS, but by applying SMC sampling it can correct for it.
FPS (Dou & Song, 2024) is also a recent method based on SMC with auxiliary variables. FPS generates a sequence of observations ${\mathbf{y}}_{1:T}$ based on a duplex diffusion process, one process at the ${\mathbf{x}}$ space and the other process at the ${\mathbf{y}}$ space. Since this method is designed for linear inverse problems only, it permits tractable Bayesian inference. Our method combines the ideas of both TDS and FPS to obtain the best of both. Namely, we use the posterior mean approximation and ${\mathbf{y}}_{1:T}$ in our SMC sampling process. As we will show, this combination can be helpful in both understanding the general semantics of an image and capturing fine details.
Several inverse sampling methods were specifically tailored for latent diffusion models. PSLD (Rout et al., 2023) extend DPS (Chung et al., 2023b) by incorporating an additional gradient update step to guide the diffusion process to sample latent representations that maintain the integrity of the decoding-encoding transformation, ensuring it remains non-lossy. STSL (Rout et al., 2024) presents a novel sampler with a tractable reverse process using an efficient second-order approximation. Comparative analysis with STSL was not feasible due to the absence of publicly available code, making replication challenging. Resample (Song et al., 2024), a contemporary method alongside PSLD, introduces a strategy for addressing general inverse problems using pretrained latent diffusion models, tackling the complexities posed by encoder and decoder nonlinearity. Resample algorithm includes hard data consistency to obtain latent variable that is consistent with the observed measurements, and then employs a resampling scheme to map the sample back onto the correct noisy data manifold and continue the reverse sampling process. Concurrent to this study Nazemi et al. (2024) proposed a particle filtering approach. Their method builds on PSLD and DPS update in the proposal distribution. Similarly to TDS (Wu et al., 2024) the connection to the labels is only through ${\mathbf{z}}_{0}$ using the approximate mean estimator. Since these methods share commonalities, we compare only to the latter in the experimental section.

### v2

@@ Abstract
In image processing, solving inverse problems is the task of finding plausible reconstructions of an image that was corrupted by some (usually known) degradation operator. Commonly, this process is done using a generative image model that can guide the reconstruction towards solutions that appear natural. The success of diffusion models over the last few years has made them a leading candidate for this task. However, the sequential nature of diffusion models makes this conditional sampling process challenging. Furthermore, since diffusion models are often defined in the latent space of an autoencoder, the encoder-decoder transformations introduce additional difficulties. To address these challenges, we suggest a novel sampling method based on sequential Monte Carlo (SMC) in the latent space of diffusion models. We name our method LD-SMC. We define a generative model for the data using additional auxiliary observations and perform posterior inference with SMC sampling based on a backward diffusion process. Empirical evaluations on ImageNet and FFHQ show the benefits of LD-SMC over competing methods in various inverse problem tasks and especially in challenging inpainting tasks.
@@ 1 Introduction
Many important signal processing tasks can be viewed as inverse problems (Song et al., 2021c; Moliner et al., 2023; Daras et al., 2024; Chung et al., 2023b; Cardoso et al., 2023). In inverse problems, the objective is to obtain a clean signal ${\mathbf{x}}\in\mathbb{R}^{n}$ from a degraded observation ${\mathbf{y}}={\mathcal{A}}({\mathbf{x}})+\bm{\mathbf{\psi}}$, where ${\mathcal{A}}$ is usually a known irreversible mapping and $\bm{\mathbf{\psi}}$ is a Gaussian noise vector. Common applications that fit this framework include image deblurring, super-resolution, inpainting, and Gaussian denoising. The broad applicability of inverse problems makes them highly significant, as they encompass numerous real-world challenges, such as those found in digital image processing (Blackledge, 2005), wireless communication (Chen et al., 2021), seismology (Virieux & Operto, 2009), medical imaging (Song et al., 2021c; Chung et al., 2023c), and astronomy (Craig & Brown, 1986).
Figure 1: LD-SMC solves inverse problem tasks in the latent space of autoencoders by utilizing auxiliary observations ${\mathbf{y}}_{1:T}$ initialized using a forward diffusion process. Then, sampling ${\mathbf{z}}_{0}$ from the posterior distribution is done based on the backward diffusion process using sequential Monte Carlo. In the figure, ${\mathcal{D}}$ and ${\mathcal{A}}$ denote the decoder and the corruption operator respectively.
@@
Recently, several studies proposed sequential Monte Carlo (SMC) (Doucet et al., 2001b; Del Moral et al., 2012) as an effective solution for this task based on pixel-space diffusion models (Cardoso et al., 2023; Trippe et al., 2023; Wu et al., 2024; Dou & Song, 2024). Specifically, Wu et al. (2024) applies the DPS approximation (Chung et al., 2023b) for $p({\mathbf{y}}|{\mathbf{x}}_{t})$ with $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ and uses SMC sampling to correct for it. Dou & Song (2024), on the other hand, connected ${\mathbf{x}}_{t}$ to ${\mathbf{y}}$ by introducing a sequence of latent variables ${\mathbf{y}}_{1:T}$ through a duplex forward diffusion process and sampling sequentially from $p({\mathbf{x}}_{t:T}|{\mathbf{y}}_{t:T})$. While this approach has shown great potential, it has two main limitations. First, it does not take into account future observations ${\mathbf{y}}_{0:t-1}$ in the sampling process. Second, and more crucially, it is limited to linear corruption models only. As such, it cannot be applied with nonlinear mappings ${\mathcal{A}}$, let alone common Latent Diffusion Models (LDMs) (Rombach et al., 2022) due to the nonlinearity of the decoder. This is a harsh restriction as many of the recent powerful and efficient models are LDMs (Esser et al., 2024).
@@
Both existing approaches have pros and cons. Using the $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ approximation can be helpful in capturing the large-scale semantics of the image, but it often lacks in capturing the small details. On the other hand, using auxiliary observations ${\mathbf{y}}_{1:T}$ can help capture finer details, but the duplex forward diffusion process is not amenable to LDMs. Here, we propose a method that combines these two approaches and strives to achieve the best of both worlds. We define a generative model for the data based on a backward diffusion process, according to which an auxiliary observation ${\mathbf{y}}_{t}$ is generated directly from ${\mathbf{z}}_{t}$. Then, to apply posterior inference over the latent diffusion variables ${\mathbf{z}}_{0:T}$, we use SMC. To obtain a tractable sampling procedure, we derive an approximate target distributions and define a novel proposal distributions for the SMC sampling process tailored for diffusion models. Hence, we name our method Latent Diffusion Sequential Monte Carlo, or more concisely LD-SMC. An illustration of our approach is shown in Figure 1. We empirically validated LD-SMC on the ImageNet (Russakovsky et al., 2015) and FFHQ (Karras et al., 2019) datasets. We found that LD-SMC outperforms or is comparable to baseline methods on image deblurring and super-resolution tasks, and can significantly improve over baseline methods on inapainting tasks, especially on the more diverse ImageNet dataset.
In this study, we make the following contributions: (1) we propose a novel method for combining auxiliary observations with latent space diffusion models; (2) we derive approximate target distributions for the SMC procedure and novel proposal distributions to perform approximate posterior sampling specifically tailored for diffusion models; (3) We theoretically show that LD-SMC is asymptotically accurate, namely that it can sample from $p_{\theta}(\mathbf{z}_{0}|\mathbf{y}_{0})$; (4) LD-SMC achieves significant improvements in perceptual quality in inpainting tasks, one of the most challenging inverse problem tasks. Our code is publicly available at https://github.com/ssi-research/LD-SMC
@@ 2 Background
Inverse Problems. In inverse problems, one would like to recover a sample ${\mathbf{x}}\in{\mathbb{R}}^{n}$ from a corrupted version of it ${\mathbf{y}}\in{\mathbb{R}}^{m}$. Usually, the corruption model that acted on ${\mathbf{x}}$ is known, but the operation is irreversible (Tarantola, 2005). For instance, restoring a high-quality image from a low-quality one. We denote the corruption operator by ${\mathcal{A}}(\cdot)$, and assume that ${\mathbf{y}}={\mathcal{A}}({\mathbf{x}})+\bm{\mathbf{\psi}}$, where $\bm{\mathbf{\psi}}\sim\mathcal{N}(0,\tau^{2}{\mathbf{I}})$ has a known standard deviation $\tau$. In a more concise way, $p({\mathbf{y}}|{\mathbf{x}})=\mathcal{N}({\mathcal{A}}({\mathbf{x}}),\tau^{2}{\mathbf{I}})$. Common examples of inverse problems are inpainting, colorization, and deblurring. In general, solving inverse problem tasks is considered an ill-posed problem with many possible solutions ${\mathbf{x}}$ with equally high $p({\mathbf{y}}|{\mathbf{x}})$ values. Given a prior distribution $p({\mathbf{x}})$ over natural images, one standard approach to solving the inverse problem is to sample from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{y}}|{\mathbf{x}})p({\mathbf{x}})$.
Diffusion Models. Owing to their high-quality generation capabilities, in recent years diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have been leveraged as priors in inverse problems (Jalal et al., 2021; Song et al., 2021d). Here, we adopt the DDIM formulation (Song et al., 2021a) for the prior model, although our approach can work with other formulations of diffusion models as well. Furthermore, since it is costly to apply the diffusion process in the pixel space, a common approach is to apply the diffusion model in the latent space given by an auto-encoder (Rombach et al., 2022). Applying diffusion models in latent space allows us to sample high-quality images while reducing the computational resources needed by the model. Hence, designing models that effectively solve inverse problems using latent diffusion models is of great importance.
Denote by ${\mathbf{z}}_{0:T}$ the random variables in the latent space. Let $\alpha_{1:T}$, $\beta_{1:T}$ be the variance schedule of the diffusion process with $\beta_{t}\coloneqq 1-\alpha_{t}$. Also, denote by $\bar{\alpha}_{t}\coloneqq\prod_{j=1}^{t}\alpha_{j}$. The DDIM sampling starts from a prior distribution at timestep $T$ set to $p({\mathbf{z}}_{T})=\mathcal{N}(0,{\mathbf{I}})$. Then, for $t>1$ the sampling is done according to $p_{\theta}({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})=\mathcal{N}({\mathbf{z}}_{t-1}|\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t},t),\mathbf{\Sigma}(t))$, where $\theta$ are the parameters of the neural network and,
@@
$\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t},t)$ $=\sqrt{\bar{\alpha}_{t-1}}\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t})+\sqrt{1-\bar{\alpha}_{t-1}-\sigma^{2}_{t}}\cdot\bm{\mathbf{\epsilon}}_{\theta}({\mathbf{z}}_{t},t).$
@@
Here we denote the approximate posterior mean of $\mathbb{E}[{\mathbf{z}}_{0}|{\mathbf{z}}_{t}]$ by $\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t})\coloneqq\frac{1}{\sqrt{\bar{\alpha}_{t}}}({\mathbf{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\mathbf{\epsilon}}_{\theta}({\mathbf{z}}_{t},t))$ (Robbins, 1956; Efron, 2011; Chung et al., 2023b). For timestep $t=1$, the mean $\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{1},1)$ is taken to be $\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{1})$. As in (Dou & Song, 2024) we fix $\sigma_{t}=\eta\cdot\sqrt{\beta_{t}\cdot\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}}$ with $\eta$ being a hyper-parameter.
@@
Sequential Monte Carlo (SMC). SMC is an important technique for sampling in probabilistic graphical models in which exact posterior inference is intractable. The SMC breaks the sampling process down to intermediate steps, allowing efficient sampling through a recursive procedure (Doucet et al., 2001b; Del Moral et al., 2012; Naesseth et al., 2019; Chopin et al., 2020).
@@
One family of probabilistic models for which SMC is especially known is state-space models (SSMs), also known as Hidden Markov Models (HMMs). In general, the following quantities must be defined in SSMs, (1) a prior distribution over the initial state $p({\mathbf{z}}_{T})$, (2) a transition distribution that defines the dynamics between states $p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})~{}\forall t<T$, and (3) a measurement model $p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})~{}\forall t<T$. The goal is to sample from the target distribution $p({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T-1})$. To do so, SMC starts by sampling $N$ particles $\{{\mathbf{z}}^{(i)}_{T}\}_{i=1}^{N}$ from the prior distribution. Then, at each step, given the previous particle set $\{{\mathbf{z}}^{(i)}_{t}\}_{i=1}^{N}$ new samples are taken from a proposal distribution ${\mathbf{z}}^{(i)}_{t-1}\sim\pi_{t-1}({\mathbf{z}}^{(i)}_{t-1}|{\mathbf{z}}^{(i)}_{t})~{}\forall i\in\{1,...,N\}$. The particles are then weighted (and possibly resampled) according to the new sequences $\{{\mathbf{z}}^{(i)}_{t-1:T}\}_{i=1}^{N}$. The proposal distribution serves as an approximation to the target distribution. Its support needs to contain the support of the target density. The weighting mechanism then corrects the approximation by assigning a weight to each particle to adjust its probability. The resampling step, if applied, aims to remove unlikely particles according to the model (Särkkä, 2013).